Solve for Y
|y+3|-4=7

Answer by rmromero(383)

To solve an equation with absolute value,
1. Simplify. Combine like terms.Add 4 both sides.
|y+3|-4 +4 = 7+4
|y+3| = 11

2. Write a disjunction, then solve the two resulting equations. y + 3 = 11 or y + 3 = -11 y + 3 - 3 = 11 - 3 or y + 3 - 3= -11 - 3 y = 8 or y = -14
Checking: |y+3|-4=7, y = 8 |8+3|-4 = 7 |11|-4 = 7 11-4 = 7 7=7 --->True
|y+3|-4=7, y = -14 |-14+3|-4=7 |-11|-4 = 7 11-4=7 7=7 --->> True
The Solution set is {8, -14}